# The Value of Information in Human-AI Decision-making

### v3

@@ Abstract
Multiple agents—including humans and AI models—are often paired on decision tasks with the expectation of achieving complementary performance, where the combined performance of both agents outperforms either one alone. However, knowing how to improve the performance of a human-AI team is often difficult without knowing more about what particular information and strategies each agent employs. We provide a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows. We demonstrate the use of the framework for model selection, empirical evaluation of human-AI performance, and explanation design. We propose a novel information-based explanation technique that adapts SHAP, a saliency-based explanation to explain information value in decision making.
@@ 1 Introduction
As the performance of artificial intelligence (AI) for domain-specific predictions improves, workflows in which human and AI models are combined to make decisions in medicine, finance, and law, among other domains. Often, statistical models can be developed that exceed the accuracy of human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954). However, whenever humans have access to additional information over the AI, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone. For example, a physician may have access to additional information that may not be captured in tabular electronic health records or other structured data (Alur et al., 2024).
@@
Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019b; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A solid understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or each individually using measures based on post hoc decision accuracy (Passi and Vorvoreanu, 2022). This approach is problematic for several reasons. First, it does not account for the best achievable performance based on the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Second, it cannot provide insight into the potential for available information to improve decision performance, making it difficult to design interventions that improve the team’s performance.
@@
In contrast, identifying information complementarities that contribute to the maximum achievable decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated information available in the environment into their judgments—provides actionable information for improving the decision pipeline. For example, the degree to which a model provides information that complements human judgments can guide AI model selection. Evidence that human experts possess decision-relevant information over AI predictions motivates further data collection to improve the AI model. Evidence that model predictions contain decision-relevant information that humans do not exploit can guide the design of explanations highlighting complementary information.
@@
We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow. In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance. Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions for a given decision problem. The intuition behind this approach is that any information that is used by the agents will eventually reveal itself through variation in their decisions. We identify the value of the information in agent (human, AI, or human-AI) decisions by offering them as a signal to a Bayesian rational decision-maker. While we focus on combinations of humans and AI agents, the methods we propose can be applied to any combination of agents.
@@
We introduce two metrics for evaluating information value in human-AI collaboration. The first—global human-complementary information value—calculates the value of a new piece of information to an agent over all of its possible realizations among all instances. The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information.
We demonstrate the framework by applying the metrics to three decision-making tasks where AI models serve as human decision-making assistants [^1] Code to replicate our experiments is available at https://osf.io/p2qzy/?view_only=ec06600d06cd4e59bb6051f992e54c08: chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019), deepfake detection (Dolhansky et al., 2020; Groh et al., 2022), and recidivism prediction (Angwin et al., 2022; Dressel and Farid, 2018). First, we demonstrate the framework’s utility in model selection by evaluating how well different AI models complement human decision-makers, showing how even among models with similar accuracy, some models strictly offer more complementary information than others across decision problems. Next, we use the framework to empirically evaluate how providing AI assistance (alongside instance-level features) helps humans exploit available information for decision-making. Lastly, we demonstrate use of the framework to design explanations by extending SHAP (Lundberg and Lee, 2017) to highlight the portion of an AI’s explanation that complements human information.
@@ 2 Related work
Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a human–AI decision pipeline.
3 Methodology
Our framework takes as input a decision problem associated with an information model and outputs the value of information of any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework. Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization drawn from the data-generating process. We also introduce a robust analysis approach to information order, which enables us to compare the agent-complementary information in signals for all possible decision problems.
@@ 3 Methodology
The choice of basic signals directly impacts how many observed samples are required to get a good estimate of the data-generating process. When high-dimensional signals such as images or text are used, it may not be computationally feasible to estimate the data-generating process. In such cases, one can preprocess the high-dimensional signals to get low-dimensional representations which (ideally) capture any important structure. These lower-dimensional signals can be defined by humans, such as when concepts are identified and then used to label the high-dimensional signals (e.g., images or parts of images). Alternatively, they can be defined algorithmically, by strategically applying dimensionality reduction to generate low-dimensional embeddings. We introduce an algorithm for identifying an “optimal” reduction model in decision problems in Appendix B. We demonstrate these two methods in Section 4 and Section 5 respectively.
@@
Our framework quantifies the value of information in a signal $V$ as the expected payoff improvement of an idealized agent who has access to $V$ in addition to some baseline information set. We suppose a rational Bayesian DM who knows the prior probability of the state and conditional distribution of signals (i.e., the data-generating process), observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff given their posterior belief. Formally, given a decision task with payoff function $S$ and an information model $\pi$, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
@@
This baseline defines the maximum expected payoff that can be achieved with no information. Bayesian decision theory quantifies the information value of $V$ by the payoff improvement of $V$ over the payoff obtained without information.
@@ 3.1 Agent-Complementary Information Value
With the above definitions, it is possible to measure the additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM. We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega}\times\mathbf{D})$ with the state and signals. The expected payoff of a Bayesian rational DM who knows $\pi$ is given by the function:
@@
Furthermore, $\operatorname{ACIV}$ can reveal complementary information between different types of agents. For instance, if we view AI predictions as $V$ and treat human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve. In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions. We demonstrate uses of $\operatorname{ACIV}$ in Section 4 and Section 5.
@@ 3.2 Instance-level Agent-complementary Information Value
Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from gaining access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions. Intuitively, this captures how much “room” there is for specific decisions to be improved through better use of the signal. To calculate instance-level information value, we rely on the performance of the Bayesian rational agent on the conditional distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b}|v)$ instead of the whole distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b})$. Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v$ on instances where $V=v$ is
$\mathrm{r}^{v,\pi,S}(v)=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v),\omega)}]$ (7)
where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v$ as defined in Equation 3. If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
$\mathrm{r}^{v,\pi,S}(v;D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v\cup d^{b}),\omega)}]$ (8)
@@
Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v$ on instances where $V=v$ as:
$\operatorname{ILIV}^{v,\pi,S}(v;D^{b})=\mathrm{r}^{v,\pi,S}(v;D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ (9)
where $\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the rational DM’s expected payoff on instances where $V=v$ with only the exisiting information encoded in agent decisions. Taking the expectation of $\operatorname{ILIV}$ over $V$ recovers the global agent-complementary information value ($\operatorname{ACIV}$), i.e.,
$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\text{\bf E}_{v\sim\pi(v)}\![{\operatorname{ILIV}^{v,\pi,S}(v;D^{b})}]$
@@ 3.2.1 ILIV-SHAP Information-based Explanation
ILIV-SHAP extends SHAP to give an explanation model on how features impact the decision-relevant information value of an individual model output.
@@
where $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ on the instances where $f(\mathbf{x})$ is realized. This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum at $\mathbf{x}^{\prime}$ where $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$, i.e., the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
@@
The explanation model offered by ILIV-SHAP is grounded in the following two properties. First, $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ are consistent with the extent to which feature $x_{i}$ contributes to the information value in $f(\mathbf{x})$. Specifically, for any two models $f$ and $f^{\prime}$ and any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$, if $x_{i}$ contributes more to the information value in $f^{\prime}$ than the information value in $f$, i.e., $\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})\geq\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})$, then $\phi^{\operatorname{ILIV}}_{i}(f^{\prime},\mathbf{x})\geq\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. This property of ILIV-SHAP follows the consistency property of SHAP (Lundberg and Lee, 2017). Second, summing up the scores $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ recovers the information value of model output, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})$. Formally, for any $f$ and any $\mathbf{x}$, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})=\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(\text{\bf E}\![{f(X)}];D^{b})+\sum_{i=1}^{m}\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. We demonstrate use of ILIV-SHAP in Section 6.
3.3 Robustness Analysis of Information Order

### v4

@@ Abstract
Multiple agents—including humans and AI models—are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually. However, knowing how to improve the performance of collaborating agents is often difficult without knowing more about what particular information and strategies each agent employs. With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows. We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information. We validate the effectiveness of the framework and ILIV-SHAP through a study of human-AI decision-making. We show that our measure of complementary information can be used to identify which AI model will best complement human decisions. We also find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP.  [^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9.
@@ 1 Introduction
As the performance of artificial intelligence (AI) models continues to improve across domains, workflows in which human experts and AI models are paired for decision-making are sought in medicine, finance, and law, among others. Statistical models can often exceed the accuracy of human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954). However, whenever humans have access to additional information over an AI model, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone. For example, a physician may have access to information that is not captured in structured health records (Alur et al., 2024).
@@
Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019b; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A rigorous understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or agents individually using measures based on post-hoc decision accuracy (Passi and Vorvoreanu, 2022). However, such approaches do not account for the best achievable performance given the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve decision performance. Consequently, it is difficult to design interventions that improve the team’s performance.
@@
If we could identify information complementarities that can be exploited to maximize the decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments—we can intervene to improve decision-making. For example, if we can identify how much complementary information over human judgments an AI model provides, we can use this knowledge to guide model selection or motivate further data collection to improve the model. Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations highlighting complementary information.
@@
We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow. In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance. Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions. While we focus on combinations of humans and AI agents, the methods we propose can be applied to any combination of agents.
@@
We introduce two metrics for evaluating information value in human-AI collaboration. The first—global human-complementary information value—calculates the value of a new piece of information to an agent over all of its possible realizations among all instances. The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information such as specific AI predictions. We derive a new explanation technique (ILIV-SHAP) that reveals how data features influence the value of complementary information for an individual prediction.
To evaluate these tools, we contribute the results of a crowdsourced between-subjects experiment in which humans make decisions with and without AI models with varying human-complementary information and with access to different explanation approaches. We find that an AI model with more human-complementary information leads to greater improvements in human-AI team performance over the human-alone baseline. We also find that adding an ILIV-SHAP explanation to a traditional SHAP explanation leads to greater improvements in human-AI performance over the human-alone baseline than only the SHAP explanation or no explanation. We demonstrate further uses of the framework for model selection, empirical evaluation of human-AI performance, and explanation design on existing datasets spanning medical diagnosis, human deepfake detection, and recidivism prediction (see Appendices E, D and F). We also include an observational study on the dataset from Vodrahalli et al. (2022) which identified the AI model with more human-complementary information helps human-AI teams achieve greater improvements in performance over the human-alone baseline in Appendix C.
@@ 2 Related work
Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a specific scenario of human–AI decision tasks.
3 Framework
Our framework takes as input a decision problem associated with an information model and outputs the value of information of any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework. The intuition behind our use of statistical decision theory is that while human decision-makers may deviate from Bayesian rational decision-making, understanding the possible value of a signal requires defining the best attainable decision performance with that signal. Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization drawn from the data-generating process. We also introduce a robust analysis approach to information order, which enables us to compare the agent-complementary information in signals for all possible decision problems.
@@ 3 Framework
The choice of basic signals directly impacts how many observed samples are required to get a good estimate of the data-generating process. When high-dimensional signals such as images or text are used, it may not be computationally feasible to estimate the data-generating process. In such cases, one can preprocess the high-dimensional signals to get low-dimensional representations which (ideally) capture any important structure. These lower-dimensional signals can be defined by humans, such as when concepts are identified and then used to label the high-dimensional signals (e.g., images or parts of images). Alternatively, they can be defined algorithmically, by strategically applying dimensionality reduction to generate low-dimensional embeddings. We introduce an algorithm for identifying a set of optimal “coarsened” signals given a set of observed samples in Appendix B.
@@
Our framework quantifies the value of information in a signal $V$ as the expected payoff improvement of an idealized agent who has access to $V$ in addition to some baseline information set. This corresponds to a rational Bayesian DM who knows the prior probability of the state and conditional distribution of signals (i.e., the data-generating process), observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff given their posterior belief. Formally, given a decision task with payoff function $S$ and an information model $\pi$, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
@@
This baseline defines the maximum expected payoff that can be achieved with no information. Bayesian decision theory quantifies the information value of $V$ as the payoff improvement of $V$ over the payoff obtained without information.
@@ 3.1 Agent-Complementary Information Value
With the above definitions, it is possible to measure the best case additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM. We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega}\times\mathbf{D})$ with the state and signals. The expected payoff of a Bayesian rational DM who knows $\pi$ is given by the function:
@@
Furthermore, $\operatorname{ACIV}$ can reveal complementary information between different types of agents. For instance, if we view AI predictions as $V$ and treat human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve. In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions. We demonstrate uses of $\operatorname{ACIV}$ in Appendix D and Appendix E.
@@ 3.2 Instance-level Agent-complementary Information Value
Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from gaining access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions. Intuitively, this captures how much “room” there is for specific decisions to be improved through better use of the signal. To calculate instance-level information value, we rely on the performance of the Bayesian rational agent on the conditional distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b}|v)$ instead of the whole distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b})$. Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v^{\prime}$ on instances with the signal realization $V=v$ is
$\mathrm{r}^{v,\pi,S}(v^{\prime})=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v^{\prime}),\omega)}]$ (7)
where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 3. If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
$\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v^{\prime}\cup d^{b}),\omega)}]$ (8)
Note that we allow $v^{\prime}\neq v$, i.e. the signal $v^{\prime}$ observed by the rational DM being different from the actual realization $v$. When $v^{\prime}\neq v$, the rational DM is misinformed, thus $\mathrm{r}^{v,\pi,S}(v^{\prime})\leq\mathrm{r}^{v,\pi,S}(v)$. $\operatorname{ILIV}$ in Definition 3.3 quantifies this improvement in payoff when the rational DM observes the actual $v$.
@@
Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v^{\prime}$ on instances where $V=v$ as:
$\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b})=\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b}).$ (9)
$\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the expected rational payoff on instances with $V=v$, where the rational DM only knows the information encoded in agent decisions $D^{b}$ without knowing any realizations of $V$. $\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v,\pi,S}(v;D^{b})\geq\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
@@ 3.2.1 ILIV-SHAP Information-based Explanation
ILIV-SHAP extends SHAP to give an explanation model of how features impact the decision-relevant information value of an individual model output.
@@
where $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ on the instances where $f(\mathbf{x})$ is realized. This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum when $g_{f}(\mathbf{x}^{\prime})$ informs $f(\mathbf{x})$, i.e., $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$ – the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
@@
The explanation model offered by ILIV-SHAP is grounded in the following two properties. First, $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ are consistent with the extent to which feature $x_{i}$ contributes to the information value in $f(\mathbf{x})$. Specifically, for any two models $f$ and $f^{\prime}$ and any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$, if $x_{i}$ contributes more to the information value in $f^{\prime}$ than the information value in $f$, i.e., $\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})\geq\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})$, then $\phi^{\operatorname{ILIV}}_{i}(f^{\prime},\mathbf{x})\geq\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. This property of ILIV-SHAP follows the consistency property of SHAP (Lundberg and Lee, 2017). Second, summing up the scores $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ recovers the information value of model output, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})$. Formally, for any $f$ and any $\mathbf{x}$, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})=\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(\text{\bf E}\![{f(X)}];D^{b})+\sum_{i=1}^{m}\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. We demonstrate use of ILIV-SHAP in Appendix F.
3.3 Robust Analysis of Information Order